susan had 4 times as many ten cent coins as fifty cent coin at first. After using 18 fifty cents coin, her number of ten cent become 6 times as many as the number of fifty cent coins.
a)How many coins does she had at first?
b)How many money does she had at first?

Answer by ankor@dixie-net.com(22740)

Let d = original no. of dimes
Let f = original no. of fifty cent coins
:
Susan had 4 times as many ten cent coins as fifty cent coin at first.
d = 4f
:
After using 18 fifty cents coin, her number of ten cent become 6 times as many as the number of fifty cent coins.
d = 6(f-18)
d = 6f - 108
:
a)How many coins does she had at first?
Replace d with 4f in the above equaation
4f = 6f - 108
108 = 6f - 4f
108 = 2f
f = 108/2
f = 54 fifty cent coins
then
4(54) = 216 ten cent coins
:
b)How many money does she had at first?
.5(54) + .10(216) =
27.00 + 21.60 = $48.60